OpenStudy (anonymous):

when it comes to translate to an inequality how would you do this then the average speed s was between 80and 90 mph would the answer be 80>s>90

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no its 80 <s < 90

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

jymmyrep is write but I would write \[80\le s \le90\]

OpenStudy (anonymous):

i would actually say \[80\leq s\leq 90\] but certainly not \[80>s>90\] which is meaningless

OpenStudy (anonymous):

thank you i see what you mean

OpenStudy (anonymous):

yes it should have been what you guys said